OpenStudy (lgbasallote):

im doing this for the latex....but it's still counted as lgbariddle :DDD \(\Large 1 + \frac{1}{1+ \frac{1}{1 + \frac{1}{1 + \frac{1}{\stackrel{\downarrow}{\infty}}}}}\)

OpenStudy (lgbasallote):

find the value :DDD

OpenStudy (lgbasallote):

@Ishaan94 good \(\LaTeX\)? hehe

OpenStudy (anonymous):

yeah lol it's the GOLDEN RATIO.

OpenStudy (lgbasallote):

golden ratio? what's that?

OpenStudy (anonymous):

\( \Large 1 + \frac{1}{1+ \frac{1}{1 + \frac{1}{1 + \frac{1}{\stackrel{\downarrow}{\infty}}}}} \) \[ x= 1+ \frac 1 x \implies x^2 - x-1 = 0 \implies x = \frac{1}{2} \left(1\pm\sqrt{5}\right) \]

OpenStudy (lgbasallote):

hah lol fast :))

OpenStudy (anonymous):

Only the + part is golden ratio.

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Golden_ratio

OpenStudy (lgbasallote):

oh i se..